OpenStudy (anonymous):

I'm 26 years old, blew off mathematics in school, and joined the workforce immediately after graduation. Now, I'm planning to go back to school for a degree in a field that (surprise!) uses that subject that "I'll never use in the real world". How realistic is it to refresh Algebra I & II and teach myself geometry, trig, and pre-calc in the next 6-12 months? I'm a reasonably intelligent person, I grasp concepts quickly, IIRC my IQ was in the mid 140's when I tested back in HS.

OpenStudy (anonymous):

PS I know this is a very generalized question and will most likely elicit responses of the 'it depends on the person' nature. For clarity, I'm inquiring more as to the difficulty in grasping the fundamental concepts of each branch.

OpenStudy (anonymous):

I think if you're headed to pure mathematics like theory based rather than applied mathematics you're understanding of the basics (Alg-PreCalc) just needs to be enough to pass through calc courses. If you're a quick learner 6 months should be all that's needed to learn the material. Now if you're headed to applied math you'll want to take the extra few months to really have a strong understanding of all the little tricks Algebra-PreCalc can give you. For example partial fractions and difficult factoring problems for the algebra all the way up to precalc's trig identities. !2 months should be plenty of time. I think it's very realistic to teach yourself the courses in 6 months... and you're pretty young. I have met tons of people in their 50s coming back to math and getting a BS in Mathematics and they are able to jump on board rather quickly. They struggle more since the way math was taught changed a bit from when they were kids. For a great site to help you review: http://patrickjmt.com/ This site is amazing for any review. He even has some humor sneaked in there for some extra fun.

OpenStudy (anonymous):

when you can do this \[\int\limits_{\pi/4}^{\pi/3}\cos^3(x)dx\] then you know you have done those subjects well, geometry is more for calc III i never understood it in highschool but have no problem with it now

OpenStudy (anonymous):

To start with, if you think this is not used in the real world, you're not well-read. Algebra requires a kind of logical thinking, the question can be figured out. It doesn't depend on learning a lot of formulas like other math. It is a challenge, like playing a game.

OpenStudy (anonymous):

The logical thinking is useful, but the math itself isn't always done out by hand or at all. Especially for business, you have programs for all the 'math' done and all you need is a head on your shoulders to go "That number is logical." Or the reverse so you can find out where you went wrong. I know what I teach in basic College Algebra, only 1 section of it is really applied straight to the the real word. That is interest rate problems. Nothing else I am required to teach applies directly to the real world. All of the logical thinking applies but not the hard math.

OpenStudy (anonymous):

Studies show that people math and engineering majors turn out to be great CEOs and businessmen.

OpenStudy (anonymous):

Thanks for all of the responses. The 'won't need in real life' comment was directed more towards the mentality of the teenager that I was. Now that I'm looking at pursuing a degree in Engineering, and have grown up a bit, I wish I'd paid a lot more attention in school. I'd definitely do high school again if I could... I wish I'd had sites like this back when I was in school.